OpenStudy (anonymous):

again. in limit! find a non zero value for the constant k that makes f(x) = tan kx/x , x < 0 3x + 2k^2 , x is bigger and equal to 0 continuous at x =0.

hartnn (hartnn):

tried it ? any ideas how to start ?

OpenStudy (anonymous):

i've tried but no solution. what i did is i substitute the x to 0 but i got 0 for each function and yet, i try to make it like this \[\frac{ \tan kx }{ x } = 3x + 2kx ^{2}\] and no solution too

hartnn (hartnn):

since the function is given continuous at x = 0 \(\Large \lim \limits_{x\to 0^-}f(x)=\lim \limits_{x\to 0^+}f(x)\)

hartnn (hartnn):

x approaches 0- means x is very near to 0, but LESS than 0 so you would use the function tan kx/x when x approaches 0- makes sense ?

hartnn (hartnn):

because f(x) is tan kx/x with x<0

hartnn (hartnn):

\(\Large \lim \limits_{x\to 0^-}\dfrac{ \tan kx }{ x }=\lim \limits_{x\to 0^+} 3x + 2kx ^{2}\) now solve those 2 limits individually

hartnn (hartnn):

the right side limit is very simple, you just plug in x=0

hartnn (hartnn):

left side limit, use tan = sin/cos \(\Large \lim \limits_{x\to 0} \dfrac{\sin x}{x} =1\)

hartnn (hartnn):

*** \(\Large \lim \limits_{x\to 0^-}\dfrac{ \tan kx }{ x }=\lim \limits_{x\to 0^+} 3x + 2k ^{2}\)

OpenStudy (anonymous):

ohhh almost forgot about that. i got the clue. let me try first. =)

OpenStudy (anonymous):

but... i got 0. hmmm?

hartnn (hartnn):

right side = 2k^2 any doubts in that ?

hartnn (hartnn):

show your try for left side limit.

OpenStudy (anonymous):

but it is actually 2kx^2

hartnn (hartnn):

you mis-typed the question ? then yes, right side =0

hartnn (hartnn):

how about left side limit ?

OpenStudy (anonymous):

got 0 too. because tan 0=0

hartnn (hartnn):

tan 0/0 = 0/0 form use tan = sin/cos

OpenStudy (anonymous):

\[\frac{ \frac{ \sin kx }{ \cos kx } }{ x }\] thats true?

hartnn (hartnn):

yesss which is just sin kx/ (x cos kx)

OpenStudy (anonymous):

the answer still zero? sorry , i'm a bit slow in limit.

hartnn (hartnn):

yes, k is 0 but the left hand limit is NOT 0

OpenStudy (anonymous):

LHL = 1?

hartnn (hartnn):